OpenStudy (jiteshmeghwal9):

Prove:- \[\Large{\cos 12^o+\cos 60^o+\cos 84^o=\cos 24^o+\cos 48^o}\]

OpenStudy (jiteshmeghwal9):

@ParthKohli

Parth (parthkohli):

\[\cos 36^{\circ} + \cos 108^{\circ} = \cos 60^{\circ}\]So write\[\cos 12^{\circ}+\cos36^{\circ}+\cos{108^{\circ}}+\cos 84^{\circ }\]\[2\cos 24^{\circ} \cos 12^{\circ} + 2\cos 96^{\circ} \cos 12^{\circ}\]\[=2\cos 12^{\circ} (\cos 24^{\circ}+\cos 96^{\circ})\]\[=2 \cos {12^{\circ}} \cdot 2 \cos 60^{\circ} \cos 36^{\circ}\]\[=2\cos12^{\circ}\cos36^{\circ}\]\[=\cos 24^{\circ} + \cos 48^{\circ}\]

OpenStudy (jiteshmeghwal9):

thanx i gt it :)